# On deformation quantization of the space of connections on a two manifold and Chern Simons Gauge Theory

### v1

@@ 1. Introduction
The construction of the star product for symplectic manifolds in general was given by deWilde and Lecomte [6] and by Fedosov [7], and for general Poisson manifolds by Kontsevich [15]. The resulting star products are often not so easy to compute.
@@
The structure of this paper is as follows. In Section 2 we repeat the computation from [18] of the Poisson bracket of functions given by polynomials in Wilson loops, giving formulas that agree with those computed by Goldman on the moduli space of flat connections. In Section 3 then recall the results of [18], and extend them to the manifold this extension is straightforward. In Section 4 we compare Chern-Simons functional integral on with deformation quantization of and show that the star product of Wilson loops arising from transversally intersecting curves, considered as functions on may be computed using the expectations defined and computed in [18]. We state this as Definition 4.4. We then prove that associativity of the star product, where defined, follows from properties of expectations in the linear functional corresponding to Chern-Simons functional integral. We close the section with an algorithm which gives an explicit formula for the star product of Wilson loops. Finally in Section 5, we briefly review the deformation quantization of Kahler manifolds arising from geometric quantization in the works of Boutet de Monvel and Guillemin [4], Bordemann, Meinrenken and Schliechenmeier [3] and Guillemin [11]. We show how applying these results to the moduli space of flat connections on a Riemann surface hints at a relation between deformation quantization of the moduli space of flat connections on a Riemann surface and the topological quantum field theory arising from geometric quantization of that space. This parallels the relation we found in this paper between deformation quantization of the space of connections and Chern-Simons gauge theory. We close by discussing the relation between the (non)-invariance of the expectations in the Chern-Simons gauge theory of [18] under the Second Reidemeister move and the (non)-invariance of the star product on under reduction. While the lack of invariance of expectations in Chern-Simons gauge theory is mysterious, the non-invariance of the star product arises from a known cause—the gauge non-invariance of the affine symplectic connection on So this hints at a possible resolution of both issues, as mentioned above.

### v2

@@ Abstract.
- 1 Introduction
- 2 The Poisson Bracket
- 3 Chern Simons gauge theory in axial gauge on
- 4 Chern Simons gauge theory in three dimensions and the star product on the space of connections in two dimensions
- 5 Some Comments and Speculations: Relation to Geometric Quantization; Quantization of Moduli spaces
- A Some comments and speculations on a coproduct structure
@@ 1. Introduction
The construction of the star product for symplectic manifolds in general was given by deWilde and Lecomte [10] and by Omori, Maeda, and Yoshioka [28] and Fedosov [12], and for general Poisson manifolds by Kontsevich [23]. The resulting star products are often not so easy to compute.
@@
The structure of this paper is as follows. In Section 2 we repeat the computation from [33] of the Poisson bracket of functions given by polynomials in Wilson loops, giving formulas that agree with those computed by Goldman on the moduli space of flat connections. In Section 3 then recall the results of [33], and extend them to the manifold this extension is straightforward. In Section 4 we compare Chern-Simons functional integral on with deformation quantization of and show that the star product of Wilson loops arising from transversally intersecting curves, considered as functions on may be computed using the expectations defined and computed in [33]. We state this as Definition 4.4. We then prove that associativity of the star product, where defined, follows from properties of expectations in the linear functional corresponding to Chern-Simons functional integral. We close the section with an algorithm which gives an explicit formula for the star product of Wilson loops. Finally in Section 5, we briefly review the deformation quantization of Kahler manifolds arising from geometric quantization in the works of Boutet de Monvel and Guillemin [6], Bordemann, Meinrenken and Schliechenmeier [5] and Guillemin [17]. We show how applying these results to the moduli space of flat connections on a Riemann surface hints at a relation between deformation quantization of the moduli space of flat connections on a Riemann surface and the topological quantum field theory arising from geometric quantization of that space. This parallels the relation we found in this paper between deformation quantization of the space of connections and Chern-Simons gauge theory. We next discuss the relation between the (non)-invariance of the expectations in the Chern-Simons gauge theory of [33] under the Second Reidemeister move and the (non)-invariance of the star product on under reduction. While the lack of invariance of expectations in Chern-Simons gauge theory is mysterious, the non-invariance of the star product arises from a known cause—the gauge non-invariance of the affine symplectic connection on So this hints at a possible resolution of both issues, as mentioned above. Finally we remark briefly on the known relation between skein modules and quantized Teichmuller space. From our point of view this provides evidence that the issues discussed above are solvable. Our results would then provide a geometric proof, based on mathematical functional integrals, of the relation between Chern Simons gauge theory and deformation quantization of moduli spaces (such as Teichmuller space). We also ask whether the coincidence we see between the results for (related to quantum invariants or the Kauffman bracket) and those for (related to Teichmuller space) would give some insight into the volume conjecture. We close with an appendix where we construct various coproducts on the functions we have studied, and ask whether any of those are related to Turaev’s coproduct.
@@ 5.1. Geometric Quantization, the star product on Kahler manifolds, and Topological Quantum Field Theory
5.3. Skein modules and quantized Teichmuller space; possible relation to the volume conjecture
A relation similar to the one we have found between quantization of the space of connections and the Kauffman bracket is the relation between skein modules and the deformation quantization of Teichmuller spaces of [14, 21]. A few references on this relation are [32, 4, 7, 8, 24, 29]. Since Teichmuller space occurs as a component of the reduction of the space of connections by the gauge group, where and since as remarked in [33] and above, the expectations in Chern Simons gauge theory are the same for and our methods should give a geometric proof of this relation using mathematical functional integrals, if we are able to find a gauge invariant symplectic connection, or at least a modification of the Chern Simons functional integral which is invariant under the second Reidemeister move, as explained in Section 5.2.
We should note that results on the uniqueness of star products give reason to expect that the desired star product, as a star product on the space of connections, is equivalent to the star product in this paper. Recall that given a star product on a symplectic manifold and a differential operator on of the form
where is another differential operator on a new star product can be defined by
for two such star products are called equivalent. Then [3, 11, 18] if any two star products are related by a combination of equivalence and a change of the parameter 272727See [18], Corollary 9.5. Since is contractible (though infinite dimensional), we may hope for a similar result here.
In a yet more speculative vein, the relation between the Kauffman bracket and quantized Teichmuller space, and more generally the coincidence of the expectations of loops in Chern Simons gauge theory for the group with those for Chern-Simons gauge theory for the groups and may be related to the volume conjecture [20, 25, 26]
Appendix A Some comments and speculations on a coproduct structure
In [2] Turaev showed that the Goldman bracket on the algebra of oriented loops on a two manifold could be supplemented with a cobracket, forming a bialgebra structure.
Something like this may be visible in our gauge theoretic construction. In particular, if we consider the case and take , the monomials in the Wilson loops might be expected to be linearly independent, and form a basis for a vector space of functions on If we take the duals of those functions to be a basis for the dual vector space then we obtain a formula similar to (2.12), giving a bracket This bracket is equivalent to a cobracket on which is the Turaev cobracket.
The bracket on can be exponentiated to give an associative product on in line with the formulas in Definition 4.4 and given more explicitly in Section 4.1. This in turn gives a coproduct on
Question A.1.
Do the product and coproduct combine to make the loops on a compact oriented two manifold into a Hopf Algebra??
If we could overcome the difficulties of gauge invariance, we might expect to obtain also a product and coproduct on the homotopy classes of loops on a compact oriented Riemann surface; equivalently, on the Wilson loops, considered as functions on In this case, there is a unit (consisting of the trivial loop) and counit (consisting of its dual), and likewise an inverse (by reversing the orientation of the loop) and antipode. We may ask whether these fit into any kind of Hopf algebra structure.
Question A.2.
Is there an analogous construction for unoriented loops?
Of course all of these speculations ignore transversality issues and other troubles, which mean that none of our products or coproducts are really well-defined. They are very far from established facts.
A.1. A geometric coproduct
There is also another coproduct construction on the algebra of loops, arising from Bott periodicity. Again, let us take
The space is a contractible space on which the group of based maps from to acts freely. The functions are -invariant functions on this space. We can therefore consider them as functions on the classifying space Note that if we are given a curve in and a point of we obtain by composition an element of taking the trace we get a complex number, and the resulting function on (or, equivalently, -invariant function on is the Wilson loop If we take these are functions on where is the space of based loops on which are homotopically trivial, by Bott periodicity. The space is a topological group (from the group structure on ) as well as an -space (from concatenation of loops). This means that the space also has two space structures. So the space of functions on acquires two coproduct structures, one of which is coassociative and one of which is coassociative ”up to homotopy”. A third construction of a coproduct arises from a map we can write directly: Identifying with the Grassmannian, consider the map arising from the map taking two -dimensional subspaces of Hilbert space to Taking the limit we obtain a map And this again gives a type of product structure on again giving a coproduct on the functions on 282828This construction on is well known (see e.g. [1]), but its application to is not, to my knowledge. Note that all these coproducts are coassociative, but not co-commutative; and they do not come in a parametrized family in any obvious way.292929It is possible to introduce an integer parameter into the third construction by modifying the map to
Question A.3.
Do any of these geometric coproducts have anything to do with the Turaev cobracket?